OpenStudy (anonymous):

solve each system by substitution. -6x+y=0 18x-3y=0

OpenStudy (australopithecus):

y = 6x sub into 18x - 3y = 0 18x - 3(6x) = 0 solve for x then sub x into one of the original equations to solve for y

OpenStudy (anonymous):

multiply the first equation by -3 then uses that equatin and add it to the 2nd equation which would eliminate the y's so then you'll have to solve for x the plug it into an original equation and find the value for y

OpenStudy (australopithecus):

Note: it doesn't matter which equation you rearrange you just need to solve for a variable in one of them and sub it into the other equation.

OpenStudy (anonymous):

ok, start off with -6x+y=0 We want to get y by itself so that we don't have to do much work. -6x+y=0 +6x +6x Add 6x to both sides. What does it look like now?

OpenStudy (australopithecus):

what you are doing in these problems is seeing where the two graphs intersect

OpenStudy (australopithecus):

just a side note on the theory behind this

OpenStudy (anonymous):

y=6x

OpenStudy (australopithecus):

I got y = 6x by rearranging: -6x+y=0

OpenStudy (anonymous):

ok, now we know what y is. So we just fit it into the second equation. 18x-3y=0 18x-3(6x)=0 Distribute, -3*6x= ?

OpenStudy (anonymous):

18x

OpenStudy (anonymous):

don't forget the negative sign, that will be dangerous. 18x-18x=0

OpenStudy (anonymous):

18x-18x=0

OpenStudy (australopithecus):

0 = 0

OpenStudy (anonymous):

add 18?

OpenStudy (anonymous):

no, they are like terms so you subtract them

OpenStudy (anonymous):

but you can have a 0x, which is just x so x=0

OpenStudy (anonymous):

Do you need help with another like this one?

OpenStudy (australopithecus):

no x doesn't equal zero

OpenStudy (australopithecus):

both equations are identical

OpenStudy (australopithecus):

18x = 18x x can be anything and they will be equal

OpenStudy (australopithecus):

18x-3y=0 18x = 3y 18x/3 = y 6x = y

OpenStudy (australopithecus):

6x = 6x

OpenStudy (australopithecus):

the lines intersect at all points

OpenStudy (anonymous):

x does equal zero, because if you fit 0 into 6x=y you get 6(0)=y so 0= y Now fit it into one of the original equations: -6x+y=0 -6(0) + (0) =0

OpenStudy (australopithecus):

the solution would be , \[(-\infty,\infty)\]

OpenStudy (anonymous):

x=0 is infinite.

OpenStudy (australopithecus):

for both variables

OpenStudy (anonymous):

austra, we are both right.

OpenStudy (australopithecus):

x = 0 is not infinite, x = 0 would give us the point (0,0) which is only one of the points of the solution. Just say that both equations are identical thus they intersect at all times

OpenStudy (australopithecus):

intersect is probably not the right word to use probably overlap would be better

OpenStudy (anonymous):

If you fit it into any of the equations it will work.

OpenStudy (anonymous):

why would you add the 18x to the other side? They're like terms, you always put like terms together.

OpenStudy (australopithecus):

@cocoa020 I dont think you understand that both equations are identical, I put them on opposite sides to show that both equations are equal

OpenStudy (anonymous):

and what does your avatar tag lead to? Is it legit?

OpenStudy (australopithecus):

x = x

OpenStudy (anonymous):

alright, ok, I'm wrong. I'll learn from it.

OpenStudy (australopithecus):

im not trying to rub it in cocoa020 :S

OpenStudy (australopithecus):

just wanted you to understand this is a tricky question

OpenStudy (anonymous):

I'm not saying you are, lol, I wish I could talk to you, because I'm trying to convey my earnest voice into text. It's coming out sarcastic.

OpenStudy (australopithecus):

well sorry for making that assumption

OpenStudy (anonymous):

No, it was a pretty good assumption considering the situation.

OpenStudy (australopithecus):

I like this problem it is one of those problems where it forces you to understand what you are doing with system of equations

OpenStudy (anonymous):

right, and I much prefer equations and formulas to angles and bisectors.